# Ke Xua,b, Fen-Fen Yanga,b, Chenggui Yuanc

### v1

@@ 3 Propagation of chaos
where the positive constant $C$ does not depend on $T$. The desired proof follows. ∎
4 Tamed-adaptive Euler-Maruyama method
4.1 Tamed coefficients and adaptive time control
For each $\Delta\in(0,1)$ define the tamed coefficients
$\sigma_{\Delta}(x,\mu)$ $:=\frac{\sigma(x,\mu)}{1+\Delta^{1/2}|\sigma(x,\mu)|(1+|x|)},$ (4.1)
$c_{\Delta}(x,\mu)$ $:=\frac{c(x,\mu)}{1+\Delta^{1/2}|c(x,\mu)|(1+|x|+|b(x,\mu)|)},$ (4.2)
$\sigma^{0}_{\Delta}(x,\mu)$ $:=\frac{\sigma^{0}(x,\mu)}{1+\Delta^{1/2}|\sigma^{0}(x,\mu)|(1+|x|)}.$ (4.3)
The adaptive time control function is taken as
$h(x,\mu)=\frac{h_{0}}{(1+|b(x,\mu)|+|\sigma(x,\mu)|+|x|^{l})^{2}+|c(x,\mu)|^{p_{0}}},$ (4.4)
with $h_{0}>0$ (we set $h_{0}=1$ in proofs), and $l,p_{0}$ as in (A3) and (A4). We will use the standard taming bounds
$|\sigma_{\Delta}(x,\mu)|(1+|x|)\leq\Delta^{-1/2},$ (4.5)
$|c_{\Delta}(x,\mu)|(1+|x|)\leq\Delta^{-1/2},$ (4.6)
$|\sigma(x,\mu)-\sigma_{\Delta}(x,\mu)|\leq C\sqrt{\Delta}\,|\sigma(x,\mu)|^{2}(1+|x|),$ (4.7)
$|\sigma^{0}(x,\mu)-\sigma^{0}_{\Delta}(x,\mu)|\leq C\sqrt{\Delta}\,|\sigma^{0}(x,\mu)|^{2}(1+|x|).$ (4.8)
4.2 Continuous-time tamed adaptive scheme with common noise
Fix $N\in\mathbb{N}$. For $i=1,\dots,N$, let $W^{i}$ be independent Brownian motions, $W^{0}$ a common Brownian motion, and $Z^{i}$ independent Lévy processes with common Lévy measure $\nu$. Define the empirical measure $\mu_{\widehat{X}^{N}_{t}}:=\frac{1}{N}\sum_{j=1}^{N}\delta_{\widehat{X}^{j,N}_{t}}$.
The continuous-time tamed adaptive Euler-Maruyama approximation $\widehat{X}^{i,N}_{t}$ is the Itô process
$\widehat{X}^{i,N}_{t}$ $=x_{0}+\int_{0}^{t}b\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}s+\int_{0}^{t}\sigma_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}W^{i}_{s}$
@@ 4.2 Continuous-time tamed adaptive scheme with common noise
$\quad+\int_{0}^{t}\sigma^{0}_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s},\mu_{\widehat{X}^{N}_{s}}\bigr{)}\,{\rm{d}}W^{0}_{s}+\int_{0}^{t}\int_{\mathbb{R}^{d}_{0}}c_{\Delta}\!\bigl{(}\widehat{X}^{i,N}_{s-},\mu_{\widehat{X}^{N}_{s-}}\bigr{)}z\,\widetilde{N}^{i}({\rm{d}}s,{\rm{d}}z),$ (4.9)
where $\widetilde{N}^{i}$ is the compensated Poisson random measure of $Z^{i}$.
Remark 4.1.
[…]
$\max_{1\leq i\leq N}\mathbb{E}\Big{[}\sup_{0\leq t\leq T}|\widehat{X}^{i,N}_{t}|^{p}\Big{]}\leq C.$
@@
The proof of this remark can be established by adapting the methodology presented in [41]. ∎

### v2

@@ 3 Propagation of chaos
for any $t\in[0,T]$, where the positive constant $C$ does not depend on time. This completes the proof. ∎
4 Tamed-Adaptive Euler-Maruyama (TAEM) Method
4.1 Discrete-time and Continuous-time TAEM Method
Set
$t_{0}:=0,\qquad X_{t_{0}}:=X_{0},$
where $X_{0}$ is the given initial random variable in (1.1). We now consider the tamed-adaptive EM scheme with step
$\Delta_{t_{k}}:=\frac{h}{1+\beta\,|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|},\qquad\beta>0,\ h\in(0,1],\qquad i.e.\qquad t_{k+1}=t_{k}+\Delta_{t_{k}},$ (4.1)
and tamed drift
$b_{\Delta_{t_{k}}}(x,\mu):=\frac{b(x,\mu)}{1+\Delta_{t_{k}}^{\alpha}|b(x,\mu)|},\qquad 0<\alpha<\frac{1}{2}.$ (4.2)
The TAEM discrete approximation is then defined by
$X_{t_{k+1}}$ $=X_{t_{k}}+b_{\Delta_{t_{k}}}\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\Delta_{t_{k}}+\sigma\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\Delta W_{t_{k}}$ (4.3)
$\quad+c\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\Delta Z_{t_{k}}+\sigma^{0}\bigl(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})\bigr)\,\Delta W^{0}_{t_{k}},$
where $\Delta W_{t_{k}}=W(t_{k+1})-W(t_{k})$, $\Delta Z_{t_{k}}=Z(t_{k+1})-Z(t_{k})$, and $\Delta W^{0}_{t_{k}}=W^{0}(t_{k+1})-W^{0}(t_{k})$. The iteration stops when $t_{k}\geq T$. Note that at each time $t_{k}$, the random variables $\Delta_{t_{k}}$ and all coefficients are $\mathcal{F}_{t_{k}}$-measurable.
The piecewise constant interpolant is defined by
$\overline{X}^{\Delta}(t):=X_{t_{k}},\qquad t\in[t_{k},t_{k+1}),$ (4.4)
and the continuous-time TAEM approximation is
$X^{\Delta}(t)$ $=X_{0}+\int_{0}^{t}b_{\Delta_{t_{s}}}\bigl(\overline{X}^{\Delta}(s),\mathcal{L}^{1}(\overline{X}^{\Delta}(s))\bigr)\,\mathrm{d}s+\int_{0}^{t}\sigma\bigl(\overline{X}^{\Delta}(s),\mathcal{L}^{1}(\overline{X}^{\Delta}(s))\bigr)\,\mathrm{d}W(s)$ (4.5)
@@ 4.1 Discrete-time and Continuous-time TAEM Method
$\quad+\int_{0}^{t}c\bigl(\overline{X}^{\Delta}(s),\mathcal{L}^{1}(\overline{X}^{\Delta}(s))\bigr)\,\mathrm{d}Z(s)+\int_{0}^{t}\sigma^{0}\bigl(\overline{X}^{\Delta}(s),\mathcal{L}^{1}(\overline{X}^{\Delta}(s))\bigr)\,\mathrm{d}W^{0}(s),$
where
$\Delta_{t_{s}}:=\Delta_{t_{k}}=\frac{h}{1+\beta|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|}\leq h,\quad{s}\in[t_{k},t_{k+1}).$
4.2 Reachability of $\Delta_{t_{k}}$ over time $T$
Theorem 4.1.
Suppose that assumptions (A1)-(A7) hold. Let $\Delta_{t_{k}}$ be given by (4.1) and define the stopping time $\tau_{T}:=\inf\left\{n\geq 0:S_{n}:=\sum_{k=0}^{n-1}\Delta_{t_{k}}\geq T\right\}.$ Then, $\mathbb{E}[\tau_{T}]<\infty$. In particular, $\tau_{T}$ is almost surely finite.
In order to prove the aforementioned theorem, the following lemmas are presented.
Lemma 4.2.
Suppose that assumptions (A1)-(A7) hold. Let $X^{\Delta}(t)$ be the continuous-time TAEM approximation given in (4.5). Then, for any real number $l\geq 2$, there exists a constant $C_{T,l}>0$, depending only on $T$, $l$ such that
$\sup_{t\leq T}\mathbb{E}[|X^{\Delta}(t)|^{l}]\leq C_{T,l}.$
In particular, since $X^{\Delta}(t_{k})=X_{t_{k}}$ for all grid points $t_{k}$, we also have
$\sup_{t_{k}\leq T}\mathbb{E}\big[|X_{t_{k}}|^{l}\big]\leq C_{T,l}.$
@@ 4.2 Reachability of $\Delta_{t_{k}}$ over time $T$
We divide the proof into several steps.
Step 1: Itô’s formula for $|X^{\Delta}(t)|^{l}$.
Let $\mu_{s}=\mathcal{L}^{1}(\overline{X}^{\Delta}(s))$ and define
$f(s)$ $:=b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\mu_{s}),\qquad g(s):=\sigma(\overline{X}^{\Delta}(s),\mu_{s}),$
$g_{0}(s)$ $:=\sigma^{0}(\overline{X}^{\Delta}(s),\mu_{s}),\qquad j(s):=c(\overline{X}^{\Delta}(s),\mu_{s}).$
Applying Itô’s formula to $|X^{\Delta}(t)|^{l}$ for $l\geq 2$
$|X^{\Delta}(t)|^{l}=|X_{0}|^{l}+l\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\,\langle X^{\Delta}(s),f(s)\rangle\,\mathrm{d}s$ (4.6)
$\quad+l\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\,\langle X^{\Delta}(s),g(s)\,\mathrm{d}W(s)\rangle+l\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\,\langle X^{\Delta}(s),g_{0}(s)\,\mathrm{d}W^{0}(s)\rangle$
$\quad+\frac{l(l-2)}{2}\int_{0}^{t}|X^{\Delta}(s)|^{l-4}\Big(|g(s)^{\top}X^{\Delta}(s)|^{2}+|g_{0}(s)^{\top}X^{\Delta}(s)|^{2}\Big)\,\mathrm{d}s$
$\quad+\frac{l}{2}\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\Big(\|g(s)\|^{2}+\|g_{0}(s)\|^{2}\Big)\,\mathrm{d}s$
$\quad+2\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\Big[|X^{\Delta}(s-)+j(s)z|^{l}-|X^{\Delta}(s-)|^{l}\Big]\,\tilde{N}(\mathrm{d}s,\mathrm{d}z)$
$\quad+\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\Big[2|X^{\Delta}(s-)+j(s)z|^{l}-2|X^{\Delta}(s-)|^{l}-l|X^{\Delta}(s-)|^{l-2}\,\langle X^{\Delta}(s-),j(s)z\rangle\Big]\,\nu(\mathrm{d}z)\,\mathrm{d}s.$
Step 2: Drift term estimation.
For the drift term $f(s):=b_{\Delta_{t_{s}}}(\overline{X}^{\Delta}(s),\mu_{s})$, by the taming property, we obtain
$|f(s)|\leq|b(\overline{X}^{\Delta}(s),\mu_{s})|.$
Using (A3) with $\overline{x}=0$ and $\overline{\mu}=\delta_{0}$, there exist constants $\tilde{L}>0$ and $\ell>0$ such that
$|b(\overline{X}^{\Delta}(s),\mu_{s})|\leq|b(0,\delta_{0})|+\tilde{L}(1+|\overline{X}^{\Delta}(s)|^{\ell})\left(|\overline{X}^{\Delta}(s)|+\mathcal{W}_{2}(\mu_{s},\delta_{0})\right).$
Since $\mathcal{W}_{2}(\mu_{s},\delta_{0})=\left(\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right)^{1/2}$, we have
$|b(\overline{X}^{\Delta}(s),\mu_{s})|$ $\leq C+\tilde{L}(1+|\overline{X}^{\Delta}(s)|^{\ell})|\overline{X}^{\Delta}(s)|+\tilde{L}(1+|\overline{X}^{\Delta}(s)|^{\ell})\left(\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right)^{1/2}$
$\leq C+\tilde{L}(1+|\overline{X}^{\Delta}(s)|^{\ell})|\overline{X}^{\Delta}(s)|+\frac{\tilde{L}}{2}(1+|\overline{X}^{\Delta}(s)|^{\ell})^{2}+\frac{\tilde{L}}{2}\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]$
$\leq C\left(1+|\overline{X}^{\Delta}(s)|^{\ell+1}+|\overline{X}^{\Delta}(s)|^{2\ell}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right).$
Now estimate the drift term, by choosing $l\geq\max\{\,2\ell,\ell+1,2\},$ we have
$\left|l\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\langle X^{\Delta}(s),f(s)\rangle\,\mathrm{d}s\right|\leq l\int_{0}^{t}|X^{\Delta}(s)|^{l-1}|f(s)|\,\mathrm{d}s$
$\leq C_{l}\int_{0}^{t}|X^{\Delta}(s)|^{l-1}\left(1+|\overline{X}^{\Delta}(s)|^{\ell+1}+|\overline{X}^{\Delta}(s)|^{2\ell}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right)\,\mathrm{d}s$
$\leq C_{l}\int_{0}^{t}\left(1+|X^{\Delta}(s)|^{l}+|\overline{X}^{\Delta}(s)|^{l}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{l}\mid\mathcal{F}^{0}]\right)\,\mathrm{d}s.$
Step 3: Diffusion terms estimation.
From Assumption (A1), we have
$|g(s)|^{2}+|g_{0}(s)|^{2}\leq L\left(1+|\overline{X}^{\Delta}(s)|^{2}+\mathcal{W}_{2}^{2}(\mu_{s},\delta_{0})\right)\leq L\left(1+|\overline{X}^{\Delta}(s)|^{2}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}]\right).$
Then we obtain
$\frac{l(l-2)}{2}\int_{0}^{t}|X^{\Delta}(s)|^{l-4}\Big(|g(s)^{\top}X^{\Delta}(s)|^{2}+|g_{0}(s)^{\top}X^{\Delta}(s)|^{2}\Big)\,\mathrm{d}s+\frac{l}{2}\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\Big(\|g(s)\|^{2}+\|g_{0}(s)\|^{2}\Big)\,\mathrm{d}s$
$\leq C_{l}\int_{0}^{t}|X^{\Delta}(s)|^{l-2}\left(|g(s)|^{2}+|g_{0}(s)|^{2}\right)\,\mathrm{d}s$
$\leq C_{l,L}\int_{0}^{t}\left[1+|X^{\Delta}(s)|^{l}+|\overline{X}^{\Delta}(s)|^{l}+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{l}\mid\mathcal{F}^{0}]\right]\,\mathrm{d}s.$
Step 4: Jump term estimation.
Write the jump contribution as $M_{J}(t)+D_{J}(t)$, where the martingale part is
$M_{J}(t)=2\int_{0}^{t}\!\!\int_{\mathbb{R}_{0}^{d}}\Big[|X^{\Delta}(s{-})+j(s)z|^{l}-|X^{\Delta}(s{-})|^{l}\Big]\,\tilde{N}(\mathrm{d}z,\mathrm{d}s),$
and the compensator is
$D_{J}(t)=\int_{0}^{t}\int_{\mathbb{R}_{0}^{d}}\Big[2|X^{\Delta}(s-)+j(s)z|^{l}-2|X^{\Delta}(s-)|^{l}-l|X^{\Delta}(s-)|^{l-2}\,\langle X^{\Delta}(s-),j(s)z\rangle\Big]\,\nu(\mathrm{d}z)\,\mathrm{d}s.$
For $l\geq 2$, by Taylor’s formula there exists $C_{l}>0$ such that for all $x,z$,
$\big||x+jz|^{l}-|x|^{l}-l|x|^{l-2}\langle x,jz\rangle\big|\;\leq\;C_{l}\big(|x|^{l-2}|jz|^{2}+|jz|^{l}\big),$
and also
$\big||x+jz|^{l}-|x|^{l}\big|\;\leq\;C_{l}\big(|x|^{l-1}|jz|^{2}+|jz|^{l}\big).$
Hence,
$|D_{J}(t)|\leq C_{l}\int_{0}^{t}\!\!\int_{\mathbb{R}_{0}^{d}}\Big(|X^{\Delta}(s)|^{l-2}|j(s)|^{2}|z|^{2}+2|j(s)|^{l}|z|^{l}+|X^{\Delta}(s)|^{l-1}|j(s)|^{2}|z|^{2}\Big)\,\nu(\mathrm{d}z)\,\mathrm{d}s.$
By Assumption (A5), we have
$|j(s)|\;\leq\;L_{3}\!\left(1+|\overline{X}^{\Delta}(s)|+\big[\mathbb{E}\!\left(|\overline{X}^{\Delta}(s)|^{2}\mid\mathcal{F}^{0}\right)\big]^{1/2}\right),$
so that $|j(s)|^{2}$ and $|j(s)|^{l}$ are bounded by $1+|\overline{X}^{\Delta}(s)|^{l}+\mathbb{E}\!\left(|\overline{X}^{\Delta}(s)|^{l}\mid\mathcal{F}^{0}\right).$ Moreover, Assumption (A4) ensures $\int_{|z|<1}|z|^{2}\nu(\mathrm{d}z)<\infty$ and $\int_{|z|\geq 1}|z|^{l}\nu(\mathrm{d}z)<\infty.$ Collecting the above, we obtain
$|D_{J}(t)|\;\leq\;C_{l,L_{3},\nu}\!\int_{0}^{t}\Big[1+|X^{\Delta}(s)|^{l}+|\overline{X}^{\Delta}(s)|^{l}+\mathbb{E}\!\left(|\overline{X}^{\Delta}(s)|^{l}\mid\mathcal{F}^{0}\right)\Big]\,\mathrm{d}s.$
Step 5: Taking expectations and Grönwall’s inequality.
Taking expectations to (4.6), all martingale terms vanish. Let
$m(t)=\sup_{0\leq u\leq t}\mathbb{E}[|X^{\Delta}(u)|^{l}],$
we obtain
$\mathbb{E}[|X^{\Delta}(t)|^{l}]\leq\mathbb{E}[|X_{0}|^{l}]+C\int_{0}^{t}\left[1+\mathbb{E}[|X^{\Delta}(s)|^{l}]+\mathbb{E}[|\overline{X}^{\Delta}(s)|^{l}]\right]\,\mathrm{d}s.$
Since $\overline{X}^{\Delta}(s)=X_{t_{k}}$ for $s\in[t_{k},t_{k+1})$, we have $\mathbb{E}[|\overline{X}^{\Delta}(s)|^{l}]\leq m(s)$ for $s\leq t$. Therefore,
$m(t)\leq\mathbb{E}[|X_{0}|^{l}]+C\int_{0}^{t}\left[1+2m(s)\right]\,\mathrm{d}s.$
By Grönwall’s inequality, we get
$m(T)\leq\left(\mathbb{E}[|X_{0}|^{l}]+CT\right)e^{2CT}=:C_{T,l}.$
This completes the proof of the lemma. ∎
Proof of Theorem 4.1.
According to Assumption (A3), we have
$|b(x,\mu)-b(0,\delta_{0})|\leq\tilde{L}(1+|x|^{\ell}+|0|^{\ell})\,(|x-0|+\mathcal{W}_{2}(\mu,\delta_{0}))=\tilde{L}(1+|x|^{\ell})(|x|+\mathcal{W}_{2}(\mu,\delta_{0})).$
Taking $x=X_{t_{k}},\ \mu=\mathcal{L}^{1}(X_{t_{k}})$, we have
$|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}})|$ $\leq C_{0}+L(1+|X_{t_{k}}|^{\ell})(|X_{t_{k}}|+\mathcal{W}_{2}(\mathcal{L}^{1}(X_{t_{k}},\delta_{0}))$
$\leq C_{0}+L(1+|X_{t_{k}}|^{\ell})\Big(|X_{t_{k}}|+(\mathbb{E}[|X_{t_{k}}|^{2}\mid\mathcal{F}^{0}])^{1/2}\Big),$
where $C_{0}=|b(0,\delta_{0})|.$ By Hölder’s inequality and Lemma 4.2, both $\mathbb{E}[(1+|X_{t_{k}}|^{\ell})|X_{t_{k}}|]$ and $\mathbb{E}[1+|X_{t_{k}}|^{\ell}\cdot(\mathbb{E}|X_{t_{k}}|^{2})^{1/2}]$ are bounded by a constant depending only on $L$ and $C_{T,\ell}$. Hence there exists $M=M(L,C_{T,\ell})$ such that,
$\mathbb{E}[|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|]\leq C_{0}+L\mathbb{E}\big[(1+|X_{t_{k}}|^{\ell})|X_{t_{k}}|\big]+L\mathbb{E}[1+|X_{t_{k}}|^{\ell}]\cdot(\mathbb{E}[|X_{t_{k}}|^{2}])^{1/2}\leq M.$
Define $f(y)=\dfrac{h}{1+\beta y}$ for $y\geq 0$. Note $f$ is convex and nonnegative on $0,\infty)$. By Jensen’s inequality,
$\mathbb{E}[\Delta_{t_{k}}]=\mathbb{E}\big[f(|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|)\big]\geq f\big(\mathbb{E}[|b(X_{t_{k}},\mathcal{L}^{1}(X_{t_{k}}))|]\big)\geq\frac{h}{1+\beta M}=:\,c>0.$
Due to $S_{\tau_{T}}-T\leq\Delta_{t_{\tau_{T}-1}}\leq h$. Fix a constant truncation integer $N$, we have
$S_{\tau_{T}\wedge N}\leq S_{\tau_{T}}\leq T+h.$ (4.7)
Taking expectations to (4.7), we have
$\mathbb{E}[S_{\tau_{T}\wedge N}]$ $=\mathbb{E}[\sum_{k=0}^{\tau_{T}\wedge N-1}\Delta_{t_{k}}]=\mathbb{E}[\sum_{k=0}^{N-1}\Delta_{t_{k}}\mathbf{1}_{\{\tau_{T}\wedge N>k\}}]=\sum_{k=0}^{N-1}\mathbb{E}[\Delta_{t_{k}}\mathbf{1}_{\{\tau_{T}>k\}}]$
$\geq\sum_{k=0}^{N-1}\mathbb{E}[\Delta_{t_{k}}]\,\mathbb{E}[\mathbf{1}_{\{\tau_{T}>k\}}]\geq c\,\mathbb{E}[\sum_{k=0}^{N-1}\mathbf{1}_{\{\tau_{T}>k\}}]=c\,\mathbb{E}[\tau_{T}\wedge N].$
Combining with $\mathbb{E}[S_{\tau_{T}\wedge N}]\leq T+h$ yields
$c\,\mathbb{E}[\tau_{T}\wedge N]\leq T+h.$
Let $N\to\infty$ and apply monotone convergence to get $c\,\mathbb{E}[\tau_{T}]\leq T+h$, so we have $\mathbb{E}[\tau_{T}]<\infty$. Therefore $\mathbb{P}(\tau_{T}=\infty)=0$ and $\tau_{T}<\infty$ almost surely. ∎
4.3 Convergence of TAEM Method under assumptions (A1)-(A7)
We now establish strong convergence of the TAEM scheme (4.3)-(4.5) under assumptions (A1)-(A7). Fix $R>0$ and define the stopping time
$\tau_{R}:=\inf\bigl\{t\geq 0:|X(t)|\geq R\ \text{or}\ |X^{\Delta}(t)|\geq R\bigr\}.$ (4.8)
Consider the stopped processes
$X^{R}(t):=X(t\wedge\tau_{R}),\qquad X^{\Delta,R}(t):=X^{\Delta}(t\wedge\tau_{R}),$
and set
$e^{R}(t):=X^{\Delta,R}(t)-X^{R}(t),\qquad e(t):=X^{\Delta}(t)-X(t).$